# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ A.2 Learning settings for figures
For this experiments, all Inception models where trained using SGD with momentum $0.9$ and learning rate $0.01$ with exponential decay of $0.95$. All models are trained for $30.000$ iterations of batches of size $200$ or until the train loss is under $0.005$. These settings are selected to ensure that the random label model converges to an interpolator. Random cropping is employed using RandomResizeCrop function of torchvision with scale $(0.8,1.0)$ and ratio $(0.9,1.1)$. For $\ell_{2}$ Regularization, we multiplicative factor is $0.01$.
@@
Figure 9.
Figure 10.
A.3 Estimating the cumulant function $J_{{\boldsymbol{\theta}}}(\lambda)$ and the rate function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$
@@
Proof  Direct consequence of the ${\cal I}_{\boldsymbol{\theta}}(a)$ being the Legendre transform of $J_{{\boldsymbol{\theta}}}(\lambda)$.
@@ B.1 The Rate Function (go back to Section 3)
Proof  First, from Assumption 1, we have that $m_{{\boldsymbol{\theta}}}\geq 0$, then $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\geq 0\ \forall({\boldsymbol{x}},{\boldsymbol{y}})\in\mathcal{X}\times\mathcal{Y}$. Then, it verifies that $\lambda(L({\boldsymbol{\theta}})-\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})))\leq\lambda L({\boldsymbol{\theta}})$. Taking exponential and expectations,
@@
By standard properties of the cumulant generating function, we have that $\frac{\partial}{\partial\lambda}J_{\boldsymbol{\theta}}(0)=0$. On the other had, also by standard properties of the cummulant (Herdegen, 2008, Lemma 1), we have that
@@
From Assumption 1 we have that $L({\boldsymbol{\theta}})<\infty$ and $m_{\boldsymbol{\theta}}\geq 0$, then we deduce that $\forall s\geq 0$ we have that ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)<\infty$.
@@
Proof  It is clear from the definition of $m_{\boldsymbol{\theta}}:=\operatorname*{ess\,inf}_{({\boldsymbol{y}},{\boldsymbol{x}})\sim\nu}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ and the rate function ${\cal I}_{\boldsymbol{\theta}}(a)$, that the limit at both sides is 00.
@@ B.2 The Rate Function Characterizes the Generalization of Interpolators (go back to Section 4)
Proof  By Chernoff’s Theorem 3.3, for a given ${\boldsymbol{\theta}}$, we have, $\mathbb{P}{}\Big{(}L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})\geq a\Big{)}\leq e^{-n{\cal I}_{\boldsymbol{\theta}}(a)}$. Naming $\delta^{\prime}=e^{-n{\cal I}_{\boldsymbol{\theta}}(a)}$ and re-arranging terms, $a={\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(-\frac{1}{n}\ln\delta^{\prime}\right)$. This allows us to rewrite the first equation as
@@
By reparametrizing the above inequality with $\delta^{\prime}=\delta k^{-p}$ we have
@@
Proof  Let $\bar{Z}_{n}:=\sqrt{n}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}}))$ and $J_{\bar{Z}_{n}},{\cal I}_{\bar{Z}_{n}}$ and ${\cal I}^{-1}_{\bar{Z}_{n}}$ denote its cumulant, rate and inverse rate function. Then, by properties of the cummulant, it verifies that $J_{\bar{Z}_{n}}(\lambda)=nJ_{{\boldsymbol{\theta}}}\big{(}\tfrac{\lambda}{\sqrt{n}}\big{)}$. Using the definition of the rate function, we can deduce that
@@
where $\lim_{t\to 0}h(t)=0$. Furthermore, it verifies that (Ellis, 2012)
@@
See 4.3
Proof  By Theorem 4.1, $L({\boldsymbol{\theta}})\leq\hat{L}(D,{\boldsymbol{\theta}})+{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$, and, by Proposition 3.1, ${\cal I}_{\boldsymbol{\theta}}(a)$ is well defined $\forall a\in[0,L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}})$. Then, by Proposition 3.2, $\forall b>0$, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(b\right)\in[0,L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}})$. In consequence, we have,
@@
The result follows from $m_{\boldsymbol{\theta}}\geq 0$, $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and rearranging terms..
See 4.4
Proof  Let $\bar{Z}_{n}:=\sqrt{n}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}}))$ and $J_{\bar{Z}_{n}},{\cal I}_{\bar{Z}_{n}}$ and ${\cal I}^{-1}_{\bar{Z}_{n}}$ denote its cumulant, rate and inverse rate function. Then, by properties of the cummulant, it verifies that $J_{\bar{Z}_{n}}(\lambda)=nJ_{{\boldsymbol{\theta}}}\big{(}\tfrac{\lambda}{\sqrt{n}}\big{)}$. Using the definition of the rate function, we can deduce that
@@
See 4.5
Proof  If ${\boldsymbol{\theta}}$ is $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$, by Definition 3, $\forall a\in(0,\beta]\quad{\cal I}_{\boldsymbol{\theta}}(a)\geq{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)$, where $\beta={\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$. Then, we have that
@@
The statement of the theorem directly derives from the above inequality.
@@ B.3 Explicit Regularization (go back to Section 5)
See 5.1
Proof  From the definitions of ${\boldsymbol{\theta}}^{\crossproduct}_{\epsilon}$ and ${\boldsymbol{\theta}}^{\star}_{\epsilon}$, given by
@@
From this, we deduce that, $L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\star}_{\epsilon})+\epsilon$. Thus, we got that $L({\boldsymbol{\theta}}^{\star}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})$ and $L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\star}_{\epsilon})+\epsilon$, finishing the proof.
See 5.2
Proof  If the loss is Lipschitz continuous with constant $M$, $\forall y,x,{\boldsymbol{\theta}}\quad\|\nabla_{{\boldsymbol{\theta}}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\|^{2}_{2}\leq M$. Then, $J_{{\boldsymbol{\theta}}}(\lambda)$ verifies
@@
See 5.3
Proof  Using Theorem 4.2, we got that $\lim_{n\rightarrow\infty}\ \sqrt{n}\,{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)=\sqrt{2\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}))\ln\tfrac{k^{p}}{\delta}}$. The proof concludes from the fact that using the exponential family,
@@
See 5.4
Proof  A second-order Taylor expansion of $J_{{\boldsymbol{\theta}}}(\lambda)$ w.r.t. to ${\boldsymbol{\theta}}$ centered around ${\boldsymbol{\theta}}_{0}$ is
@@
Evaluating the expression on $s=\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}$ and ${\boldsymbol{\theta}}=\mathbf{0}$ concludes the proof.
See 5.5
Proof  Using Chafaï (2004)’s Corollary 2.1 on $\phi(f)=-\ln f$ and $f_{\boldsymbol{\theta}}(y,{\boldsymbol{x}})=e^{-\lambda\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})}$; we get that
@@ B.4 Invariances (go back to Section 6)
See 6.1
Proof  According to Assumption 2, the targets ${\boldsymbol{y}}$ are independent of the input ${\boldsymbol{x}}_{t}$ given ${\boldsymbol{x}}_{t-1}$, that is, ${\boldsymbol{y}}\perp{\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{t-1}$. This is a result of the fact that ${\boldsymbol{x}}_{t-1}$ d-connects ${\boldsymbol{x}}_{t}$ and ${\boldsymbol{y}}$ in the graph representation of Assumption 2. The result follows then by the data processing inequality.
See 6.2
Proof  By definition we have $L^{\nu_{t+1}}({\boldsymbol{\theta}})={\mathbb{E}}_{\nu_{t+1}}[\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t+1}),{\boldsymbol{\theta}})]={\mathbb{E}}_{\nu_{t}}{\mathbb{E}}_{g_{t+1}}[\ell({\boldsymbol{y}},g_{t+1}({\boldsymbol{x}}_{t}),{\boldsymbol{\theta}})]$. As $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ is convex in ${\boldsymbol{x}}$, the expectation under $g_{t+1}$ can be moved inside $\ell$:
@@
Given that ${\mathbb{E}}_{g_{t+1}}[g_{t+1}({\boldsymbol{x}})]={\boldsymbol{x}}$, we got $L^{\nu_{t+1}}({\boldsymbol{\theta}})\geq{\mathbb{E}}_{\nu_{t}}[\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t},{\boldsymbol{\theta}})]=L^{\nu_{t}}({\boldsymbol{\theta}})$.
See 6.3
Proof  First of all we will show that $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)\geq J^{\nu_{t}}_{{\boldsymbol{\theta}}}(\lambda)$; from there, the result is immediate due to definition of the inverse rate. By definition, we got $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)=\lambda L^{\nu_{t+1}}({\boldsymbol{\theta}})+\ln{\mathbb{E}}_{\nu_{t+1}}[e^{-\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t+1},{\boldsymbol{\theta}})}]$, where $\nu_{t+1}$ can be expanded leading to
@@
From this inequality and by considering standard properties of the Legendre transform, we have $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)\geq J^{\nu_{t}}_{{\boldsymbol{\theta}}}(\lambda)\implies{\cal I}^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(a)\leq{\cal I}^{\nu_{t}}_{{\boldsymbol{\theta}}}(a)$.
See 6.4
Proof  It immediately follows from the definition of model invariant (Definition 4) and the definitions of $L^{\nu_{t+1}}({\boldsymbol{\theta}})$, $L^{\nu_{t}}({\boldsymbol{\theta}})$, ${\cal I}^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(a)$ and ${\cal I}^{\nu_{t}}_{{\boldsymbol{\theta}}}$.
See 6.5
Proof  Under Assumption 2, it is clear that
@@
From this point, the inequality regarding the inverse rate is clear from the definition.
See 6.6
Proof  Under Assumption 2, it is clear that
@@
See 6.7
Proof  On one hand, by definition of $\ell_{G}$ it verifies that
@@
See 6.8
Proof  First of all, by definition we got that $m_{\boldsymbol{\theta}}=\operatorname*{ess\,inf}_{({\boldsymbol{x}},{\boldsymbol{y}})\sim\nu_{T}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$. Then, it verifies that
@@
As a result, the empirical loss $\hat{L}^{\ell}(D_{0},{\boldsymbol{\theta}})$ is equal to $m_{\boldsymbol{\theta}}$ too.

### v4

@@ A.1 Learning Settings for Figures
For this experiments, all Inception models where trained using SGD with momentum $0.9$ and learning rate $0.01$ with exponential decay of $0.95$. All models are trained for $30.000$ iterations of batches of size $200$ or until the train loss is under $0.005$. These settings are selected to ensure that the random label model converges to an interpolator. Random cropping is employed using RandomResizeCrop function of torchvision with scale $(0.8,1.0)$ and ratio $(0.9,1.1)$. For $\ell_{2}$ regularization, the multiplicative factor is $0.01$.
@@
A.2 Estimating the Cumulant Function $J_{{\boldsymbol{\theta}}}(\lambda)$ and the Rate Function ${\cal I}_{\boldsymbol{\theta}}(\cdot)$
@@
Proof.
Direct consequence of the ${\cal I}_{\boldsymbol{\theta}}(a)$ being the Legendre transform of $J_{{\boldsymbol{\theta}}}(\lambda)$. ∎
@@ B.1 The Rate Function (go back to Section 3)
Proof.
First, from Assumption 1, we have that $m_{{\boldsymbol{\theta}}}\geq 0$, then $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\geq 0\ \forall({\boldsymbol{x}},{\boldsymbol{y}})\in\mathcal{X}\times\mathcal{Y}$. Then, it verifies that $\lambda(L({\boldsymbol{\theta}})-\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})))\leq\lambda L({\boldsymbol{\theta}})$. Taking exponential and expectations,
@@
By standard properties of the cumulant generating function, we have that $\frac{\partial}{\partial\lambda}J_{\boldsymbol{\theta}}(0)=0$. On the other had, also by standard properties of the cummulant (Herdegen,, 2008, Lemma 1), we have that
@@
From Assumption 1 we have that $L({\boldsymbol{\theta}})<\infty$ and $m_{\boldsymbol{\theta}}\geq 0$, then we deduce that $\forall s\geq 0$ we have that ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(s\right)<\infty$. ∎
@@
Proof.
It is clear from the definition of $m_{\boldsymbol{\theta}}:=\operatorname*{ess\,inf}_{({\boldsymbol{y}},{\boldsymbol{x}})\sim\nu}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ and the rate function ${\cal I}_{\boldsymbol{\theta}}(a)$, that the limit at both sides is 00. ∎
@@ B.2 The Rate Function Characterizes the Generalization of Interpolators (go back to Section 4)
Proof.
By Chernoff’s Theorem 3.3, for a given ${\boldsymbol{\theta}}$, we have, $\mathbb{P}{}\Big{(}L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}})\geq a\Big{)}\leq e^{-n{\cal I}_{\boldsymbol{\theta}}(a)}$. Naming $\delta^{\prime}=e^{-n{\cal I}_{\boldsymbol{\theta}}(a)}$ and re-arranging terms, $a={\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(-\frac{1}{n}\ln\delta^{\prime}\right)$. This allows us to rewrite the first equation as
@@
By reparameterizing the above inequality with $\delta^{\prime}=\delta k^{-p}$ we have
@@
Proof.
Thet proof follows the same approach as Theorem 4.1, with the union bound applied to the models in $\bar{{\boldsymbol{\Theta}}}$ rather than directly to the models in ${\boldsymbol{\Theta}}$. This adjustment is valid because, when applying the union bound, we only need to consider the total number of distinct random variables; in this case, $\hat{L}(D,{\boldsymbol{\theta}})$ with $D\sim\nu^{n}$. Models in ${\boldsymbol{\Theta}}$ that are not in $\bar{{\boldsymbol{\Theta}}}$ define random variables that are duplicated by definition and therefore do not need to be accounted for in the application of the union bound. ∎
See 4.3
Proof.
Let $\bar{Z}_{n}:=\sqrt{n}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}}))$ and $J_{\bar{Z}_{n}},{\cal I}_{\bar{Z}_{n}}$ and ${\cal I}^{-1}_{\bar{Z}_{n}}$ denote its cumulant, rate and inverse rate function. Then, by properties of the cummulant, it verifies that $J_{\bar{Z}_{n}}(\lambda)=nJ_{{\boldsymbol{\theta}}}\big{(}\tfrac{\lambda}{\sqrt{n}}\big{)}$. Using the definition of the rate function, we can deduce that
@@
where $\lim_{t\to 0}h(t)=0$. Furthermore, it verifies that (Ellis,, 2012)
@@
See 4.4
Proof.
By Theorem 4.1, $L({\boldsymbol{\theta}})\leq\hat{L}(D,{\boldsymbol{\theta}})+{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\textstyle\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$, and, by Proposition 3.1, ${\cal I}_{\boldsymbol{\theta}}(a)$ is well defined $\forall a\in[0,L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}})$. Then, by Proposition 3.2, $\forall b>0$, ${\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(b\right)\in[0,L({\boldsymbol{\theta}})-m_{\boldsymbol{\theta}})$. In consequence, we have,
@@
The result follows from $m_{\boldsymbol{\theta}}\geq 0$, $\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon$ and rearranging terms.. ∎
See 4.5
Proof.
Let $\bar{Z}_{n}:=\sqrt{n}(L({\boldsymbol{\theta}})-\hat{L}(D,{\boldsymbol{\theta}}))$ and $J_{\bar{Z}_{n}},{\cal I}_{\bar{Z}_{n}}$ and ${\cal I}^{-1}_{\bar{Z}_{n}}$ denote its cumulant, rate and inverse rate function. Then, by properties of the cummulant, it verifies that $J_{\bar{Z}_{n}}(\lambda)=nJ_{{\boldsymbol{\theta}}}\big{(}\tfrac{\lambda}{\sqrt{n}}\big{)}$. Using the definition of the rate function, we can deduce that
@@
See 4.6
Proof.
If ${\boldsymbol{\theta}}$ is $\beta$-smoother than ${\boldsymbol{\theta}}^{\prime}$, by Definition 3, $\forall a\in(0,\beta]\quad{\cal I}_{\boldsymbol{\theta}}(a)\geq{\cal I}_{{\boldsymbol{\theta}}^{\prime}}(a)$, where $\beta={\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)$. Then, we have that
@@
The statement of the theorem directly derives from the above inequality. ∎
@@ B.3 Explicit Regularization (go back to Section 6)
See 6.1
Proof.
From the definitions of ${\boldsymbol{\theta}}^{\crossproduct}_{\epsilon}$ and ${\boldsymbol{\theta}}^{\star}_{\epsilon}$, given by
@@
From this, we deduce that, $L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\star}_{\epsilon})+\epsilon$. Thus, we got that $L({\boldsymbol{\theta}}^{\star}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})$ and $L({\boldsymbol{\theta}}^{\crossproduct}_{\epsilon})\leq L({\boldsymbol{\theta}}^{\star}_{\epsilon})+\epsilon$, finishing the proof. ∎
See 6.2
Proof.
If the loss is Lipschitz continuous with constant $M$, $\forall y,x,{\boldsymbol{\theta}}\quad\|\nabla_{{\boldsymbol{\theta}}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})\|^{2}_{2}\leq M$. Then, $J_{{\boldsymbol{\theta}}}(\lambda)$ verifies
@@
See 6.3
Proof.
Using Theorem 4.3, we got that $\lim_{n\rightarrow\infty}\ \sqrt{n}\,{\cal I}^{-1}_{{\boldsymbol{\theta}}}\left(\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}\right)=\sqrt{2\mathbb{V}_{\nu}(\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}}))\ln\tfrac{k^{p}}{\delta}}$. The proof concludes from the fact that using the exponential family,
@@
See 6.4
Proof.
A second-order Taylor expansion of $J_{{\boldsymbol{\theta}}}(\lambda)$ w.r.t. to ${\boldsymbol{\theta}}$ centered around ${\boldsymbol{\theta}}_{0}$ is
@@
Evaluating the expression on $s=\tfrac{1}{n}\ln\tfrac{k^{p}}{\delta}$ and ${\boldsymbol{\theta}}=\mathbf{0}$ concludes the proof. ∎
See 6.5
Proof.
Using Chafaï, (2004)’s Corollary 2.1 on $\phi(f)=-\ln f$ and $f_{\boldsymbol{\theta}}(y,{\boldsymbol{x}})=e^{-\lambda\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})}$; we get that
@@ B.4 Invariances (go back to Section 7)
See 7.1
Proof.
According to Assumption 2, the targets ${\boldsymbol{y}}$ are independent of the input ${\boldsymbol{x}}_{t}$ given ${\boldsymbol{x}}_{t-1}$, that is, ${\boldsymbol{y}}\perp{\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{t-1}$. This is a result of the fact that ${\boldsymbol{x}}_{t-1}$ d-connects ${\boldsymbol{x}}_{t}$ and ${\boldsymbol{y}}$ in the graph representation of Assumption 2. The result follows then by the data processing inequality. ∎
See 7.2
Proof.
By definition we have $L^{\nu_{t+1}}({\boldsymbol{\theta}})={\mathbb{E}}_{\nu_{t+1}}[\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t+1}),{\boldsymbol{\theta}})]={\mathbb{E}}_{\nu_{t}}{\mathbb{E}}_{g_{t+1}}[\ell({\boldsymbol{y}},g_{t+1}({\boldsymbol{x}}_{t}),{\boldsymbol{\theta}})]$. As $\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$ is convex in ${\boldsymbol{x}}$, the expectation under $g_{t+1}$ can be moved inside $\ell$:
@@
Given that ${\mathbb{E}}_{g_{t+1}}[g_{t+1}({\boldsymbol{x}})]={\boldsymbol{x}}$, we got $L^{\nu_{t+1}}({\boldsymbol{\theta}})\geq{\mathbb{E}}_{\nu_{t}}[\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t},{\boldsymbol{\theta}})]=L^{\nu_{t}}({\boldsymbol{\theta}})$. ∎
See 7.3
Proof.
First of all we will show that $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)\geq J^{\nu_{t}}_{{\boldsymbol{\theta}}}(\lambda)$; from there, the result is immediate due to definition of the inverse rate. By definition, we got $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)=\lambda L^{\nu_{t+1}}({\boldsymbol{\theta}})+\ln{\mathbb{E}}_{\nu_{t+1}}[e^{-\ell({\boldsymbol{y}},{\boldsymbol{x}}_{t+1},{\boldsymbol{\theta}})}]$, where $\nu_{t+1}$ can be expanded leading to
@@
From this inequality and by considering standard properties of the Legendre transform, we have $J^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(\lambda)\geq J^{\nu_{t}}_{{\boldsymbol{\theta}}}(\lambda)\implies{\cal I}^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(a)\leq{\cal I}^{\nu_{t}}_{{\boldsymbol{\theta}}}(a)$. ∎
See 7.4
Proof.
It immediately follows from the definition of model invariant (Definition 4) and the definitions of $L^{\nu_{t+1}}({\boldsymbol{\theta}})$, $L^{\nu_{t}}({\boldsymbol{\theta}})$, ${\cal I}^{\nu_{t+1}}_{{\boldsymbol{\theta}}}(a)$ and ${\cal I}^{\nu_{t}}_{{\boldsymbol{\theta}}}$. ∎
See 7.5
Proof.
Under Assumption 2, it is clear that
@@
From this point, the inequality regarding the inverse rate is clear from the definition. ∎
See 7.6
Proof.
Under Assumption 2, it is clear that
@@
See 7.7
Proof.
On one hand, by definition of $\ell_{G}$ it verifies that
@@
See 7.8
Proof.
First of all, by definition we got that $m_{\boldsymbol{\theta}}=\operatorname*{ess\,inf}_{({\boldsymbol{x}},{\boldsymbol{y}})\sim\nu_{T}}\ell({\boldsymbol{y}},{\boldsymbol{x}},{\boldsymbol{\theta}})$. Then, it verifies that
@@
As a result, the empirical loss $\hat{L}^{\ell}(D_{0},{\boldsymbol{\theta}})$ is equal to $m_{\boldsymbol{\theta}}$ too. ∎